OpenStudy (anonymous):

5/9 x 27= ? please help

OpenStudy (anonymous):

do you know how to multiply fractions together?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\(\Huge \frac{a}{c} * \frac{b}{d} = \frac{a*b}{c*d}\)

OpenStudy (anonymous):

if you're trying to multiply 2 fractions together you multiply the numerators together and you multiply the denominators togeter look at what smooth as written

OpenStudy (anonymous):

It helps to express both numbers as a fraction. In this case, rewrite 27 as \(\Large \frac{27}{1}\)

OpenStudy (anonymous):

15

OpenStudy (anonymous):

5/9x27/1=15